Consider the following three bond quotes: a Treasury note quoted at 96.719, a corporate bond quoted at 103.70, and a municipal bond quoted at 102.35. If the Treasury and corporate bonds have a par value of $1,000 and the municipal bond has a par value of $5,000, what is the price of these three bonds in dollars? (Do not round intermediate calculations and round your final answers to 2 decimal places.)

Treasury bond=

Corporate bond=

Municipal bond=
Answer #1

price in dollars = quote*par value/100

Treasury bond=96.719*1000/100=967.19

Corporate bond=103.7*1000/100=1037

Municipal bond=102.35*5000/100=5117.5